Problem 21.11

A flat, square coil with 17.0 turns has sides of length 0.140 m . The coil rotates in a magnetic field of 2.00×10−2 T .

Part A

What is the angular velocity of the coil if the maximum emf produced is 20.0 mV ?

ω = 3.00   rad/s  

SubmitHintsMy AnswersGive UpReview Part

Correct

Part B

What is the average emf at the angular velocity found in part A?

〈E〉 =
Answer #1

length of the coil = l = 0.14 m

area of cross section the coil = A = l^2 = 0.14^2

magnetic field B = 2*10^-2 T

magnetic flux phi = B*A*coswt

induced emf = rate of change in flux

emf = (d/dt)*B*A*coswt

emf = B*A*sinwt*w

maximum emf = B*A*w


given maximum emf = 20*10^-3 v

w = (20*10^-3)/(0.14^2*2*10^-2)


w = 51.02 rad/s


++++++


partB


average value of sinwt = 1/2

average emf = maxx emf/2

average = 20/2 mv


average = 10 mV <<<-------------answer